log(x^3)+log(x^4)/log(80x)=3
Answer by ewatrrr(24785)

log(x^7) = 3log(80x)
log(x^7) = log(80x)^3
x^7 = 80^3x^3
x^4 = 80^3
x = ± root(4, 512,000) =± 26.7496 rounded